OpenStudy (the_rebel_outlaw):

The volume of a gas is 605 liters at 27.0°C. The new temperature is -3.0°C. What is the new volume? (Remember to convert temperatures to the Kelvin scale.)

OpenStudy (dumbcow):

V = kT Find k by plugging in given V,T \[k = \frac{605}{27+273}\] Now find new Volume given new T \[V = \frac{605}{300}(-3+273)\] \[V = 544.5\]